You deposit $3000 each year into an account earning 3% interest compunded annually. How much will you have in the account in 20 years?
Answer #1

Given, P = $3000

r = 3% interest compounded annually

t = 20 years

We know that in t years the amount will be = P(1+\frac{r}{100})\frac{(1+\frac{r}{100})^{t}-1}{(1+\frac{r}{100})-1}

Then, in 20 years, the amount will be = $[3000(1+\frac{3}{100})\frac{(1+\frac{3}{100})^{20}-1}{(1+\frac{3}{100})-1}]

= $[3090*\frac{1.03^{20}-1}{0.03}]

\approx $83029.46

Therefore, i will have $83029.46 in the account in 20 years.
Answer #2

You deposit $3000 each year into an account earning 3% interest compounded annually. How much will you have in the account in 15 years?

source: math
answered by: TD3215